Home‎ > ‎Reports‎ > ‎Crew 108‎ > ‎

day05


Crew #108 Log Book for December 07, 2011

Photos of the day





Commander's Check-In Report

Written by :

Charlotte Poupon, Commander


Crew Physical Status:

One crew member stayed the night in a motel in Hanksville due to tiredness and possible recurrent kidney infection.

The commander advised the film crew that due to hard living conditions in confinement, and the food study experiment running, the sick crew member might not complete the rest of the simulation. The crew was given hospital and pharmacy phone numbers. They will let us know about the health of the crew member and come back tomorrow around 7:30.

Time Departed/Returned from EVA:

EVA # 9 / 7:40 to 9:45
EVA # 10 / 15:30 to 17:30


Brief Narrative of Field Mission Results:

The cinema crew is very satisfied with their shots. They will come back tomorrow for their last day of shooting.


EVA Data/Interpretations:

Not applicable
Engineering/Hab Maintenance:
Not applicable


Report Transmission Schedule:

All reports by 22:00.


Plans for Tomorrow:

Go on filming with the crew If we have time, Light measurement installation (luxmetre) will be put up.


Inventory: 

Not applicable


Support Requested: 

Not applicable

Commander's Report

Date and time:

7 December 2001


Written by : 

Charlotte Poupon, Commander

Text:

The day was pretty calm and the sim didn't have any scientific purpose since we are playing for the filming crew. When half of the crew members were on EVA, the others stayed at the Hab to do engineering work like fixing suits. This will be documented in the Engineering Report later on.

One crew member's participation in the mission is unclear. It has to be settled by tomorrow since the filming crew is flying back to Europe on the day after
tomorrow. In case the crew member will not be returning to the mission, crew members agreed to raise the level of simulation. The Commander will assume the crew member's shifts in the daily routines.

Thanks to an email from Peter Detterline we were told that the Observatory is
open, so we'll be glad to follow all the procedures to operate it very soon.

The general mood is really ok. We took our first showers… seems to be good, after drying quickly !

Engineering Report

Date and time: 

20:30, 7 December 2011


Written by: 

Ashley Dale & Mike Lotto, Crew 108, Engineers


Generator/Electricity:
Honey (Diesel Generator) (Used? Yes/No): Nominal
Wendy (Gas Generator) (Used? Yes/No): No

Battery Charge Level: Unknown
Notes/Comments (include how many generator runs, approximately when, and any times you needed to turn the generator on/off): None

Fuel Status: N/A
ATV Oil Quantity (# of Quarts in storage): Unknown
Car Oil Quantity (# of Quarts in storage): Unknown

Water Status:
Outside Potable Water Tank Level (inches from bottom): 0.5"
Trailer Potable Water Tank Level (inches from top): 13"
Hab Potable Water Tank Level (inches from bottom): Full 
 
Observatory:
Telescope Used? (Yes/No): No
Observatory Notes/Comments: N/A 

Transportation:
Hab Car (NOT RENTAL)
Used (yes/no): No
Oil Checked (yes/no): No
Tire Status: Nominal
Odometer at end of day: Unknown
Notes/Comments on Hab Car: 
 
ATV 1:
Used: Yes
Oil Checked: No
Fuel Consumed: 0.2 gallons
Tire Status: Good
Comments: Replace oil tomorrow 
 
ATV 2:
Used: Yes
Oil Checked: No
Fuel Consumed: 0.2 gallons
Tire Status: Good
Comments: Replace oil tomorrow 
 
ATV 3:
Used: Yes
Oil Checked: No
Fuel Consumed: 0.2 gallons
Tire Status: Good
Comments: Replace oil tomorrow

Heating and Ventilation:
Thermostat setting upstairs (Farenheit): 65
Actual temperature upstairs (Farenheit): 65
Thermostat setting downstairs (Farenheit): 65
Actual temperature downstairs (Farenheit): 
 
Computers/Networking Infrastructure: 
Webcams functioning fine but not uploading 

Notes/comments on internet/computers?: 
After resetting Hab laptop (Acer), internet functionality improved.

General Engineering Notes/Comments: 
EVA suits 1, 2, 3, 4 had maintenance checks and repairs, need to check fan in toilet (rattling too much), need to find oil for ATVs/HabCar

Journalist Report


Date and Time: 

17:30, 7 December 2011


Written by: 

Usha Lingappa, Crew 108, Biologist


This was our first non-cooking day of the food study, and it went smoothly. The day centered around continued work on the film. Ashley, Michael, Usha, and Alicia went on a morning EVA with the film crew. One crew member accompanied the film crew and left before lunch to scout out locations for future shots out of sim. When they hadn't yet returned by 3pm, we began to worry about them. 

The crew eventually returned without the crew member, reporting that wasn't feeling well, had remained in Hanksville, and was considering going to the hospital later tonight. Lotto, Michael, and Charlotte went on an afternoon EVA with the film crew to continue shooting. The continued participation of the sick crew member in the mission is unclear. We plan to intensify the level of simulation for the remainder of the mission once the filming is complete.

EVA 9 Report

Date and time: 

13:30,  7 December 2011


Written by: 

Ashley Dale, Crew 108, Engineer


Time: 

07:40 – 09:45


Crew Members: 

Ashley Dale, Usha Lingappa, Michael LeClair, Alicia Framis


Site Location: 

North of the Thieltges Flats and south of Schubert Planitia


Transit Mode:

New Blue

Objectives:

Film crew for documentary required a few particular shots to be taken. Michael explored a few new locations regards geology.


Results: 

Film crew got the shots they were after. Michael explored a few new locations but found no places of significant interest.


Lessons Learned:

N/A


Other